# Di-NeRF: Distributed NeRF for Collaborative Learning with Unknown Relative Poses

### v1

@@ IV-C Trajectory Overlap Analysis
In this experiment, the sensitivity of the convergence of the relative poses of the robots with respect to the overlaps between the views of the robots is analyzed. For this experiment, two robots are considered. There is no common images between the robots, and the overlaps are only in the trajectories. Fig. 8 shows the top-down view of the overlaps, ranging from 5% to 40%, i.e. [5, 10, 20, 30, 40]%. The overlaps are determined using $x$ and $y$ coordinates of camera poses, assuming the poses are known in a coordinate frame. To divide the views among $N$ robots, first, the $x-y$ plane containing the desired object (the Chair, in this case) is divided into $N$ sectors, centered on the object. This will generate a 0% overlap for $N$ robots, which then is extended to increase the overlap, as shown in Fig. 8.
For two robots, once the overlaps are determined, the poses of robot $R^{2}$ are manipulated to create a relative displacement with respect to $R^{1}$, by translating them for 100 cm along each axis and rotating them for ${}^{\circ}$ around each axis. Once the data is created, optimization is performed, assuming an initial estimate for the relative pose with 60 cm for each translation component and ${}^{\circ}$ for each rotation.
Figure 8: Segmenting the Chair sequence with different overlaps.
Figure 9: Optimizing the relative poses for two robots with varying overlaps. The average of translation and rotation estimates are plotted. The desired translation/rotation is marked by GT. Large overlaps improve the convergence.
@@ IV-D Waymo dataset - Unbounded Scenes
The San Francisco Mission Bay Dataset [10] is collected in an urban landscape with dynamic objects and reflective surfaces. It records 12,000 images with an array of 12 cameras, with different viewpoints, on a car for 100 seconds, over a distance of 1.08 km. For this experiment, a segment of the data, including 233 images from one camera (one that provides a complete surround view from the roof of the car), measuring approximately 286 meters was selected and shown in Fig. 10. The sequence was divided into six segments to resemble a multi-agent setup. The primary objective of this experiment is not to surpass the state-of-the-art of Block-NeRF [10]. The goal here is to achieve comparable results between a distributed and a centralized system. In Fig. 11 the rendering result for this setup is shown for centralized and Di-NeRF training. All robots are able to render a view that they have never seen directly. The average values for the PSNR and SSIM metric are 25.10 and 0.814 over 6 robots and 25.21 and 0.825 for the centralized setup, In Table IV, the values for each robot are presented.
@@
Figure 10: A sequence of the Mission Bay Dataset trajectories is divided into six segments, i.e. robots, and used to evaluate Di-NeRF.
@@
Figure 11: Rendered images from Di-NeRF on Mission Bay Dataset, all six robots demonstrated the capability to render similar images with quality comparable to that achieved through centralized processing.
Table IV: Di-NeRF and centralized image metric performances for Mission Bay Dataset for 6 robots, Lower/higher better, is indicated by $\downarrow$/$\uparrow$ respectively. PSNR and SSIM are reported for each robot and relative pose optimization accuracy for all robots except robot $R^{1}$ which is considered as the global coordinate.
R#1 R#2 R#3 R#4 R#5 R#6 Cntr.
@@
SSIM $\uparrow$ 0.814 0.816 0.814 0.817 0.815 0.809 0.825
$\delta R^{\circ}\downarrow$ 0.85 0.81 0.81 0.87 0.88 -
$\delta t(cm)\downarrow$ 3.1 1.9 2.6 2.9 3.1 -
IV-E Technical Details of the Experiments
In this work, three datasets are used to evaluate Di-NeRF: The synthetic dataset, the Tanks and Temples, and the Waymo dataset. Consistency was maintained in these experiments concerning several parameters:
Batch Size: For all datasets is 2048.
Learning Rate: For all experiments is 0.01.
Regularization Parameter: The value of $\rho$ in (9) is 0.001, which is the weight for the quadratic term and is the step size in the gradient ascent for the dual variable optimization.
Communication Rounds: The number of steps before a communication round is 10 for all robots.
NeRF Network Configuration: The network $F_{\theta}$ has 8 layers, each with 256 channels. The hashmap size is $2^{19}$.
Dataset Specifications: Image size for the Synthetic Dataset is 800$\times$800 pixels, for Tanks and Temples is 1920$\times$1080 pixels, and for the Waymo Dataset is 1217$\times$1096 pixels. The rendering time for a fully connected graph is summarized in table V.
Table V: Di-NeRF and centralized training time comparison.
Dataset Di-NeRF (sec/itr) Centralized (sec/itr) N_Robots
Synthetic Dataset 4.27 0.94 2
Tank and Temple 4.78 0.93 5
Mission Bay 4.73 0.95 6
@@ V Conclusion and Future Work
We introduced the Di-NeRF algorithm, the first fully distributed NeRF used for multi-robot systems and 3D scenes, facilitating high-performance distributed training of Neural Radiance Fields (NeRF). We demonstrate its versatility across synthetic and real datasets. Di-NeRF harnesses the advantages of NeRF in distributed learning, relying solely on RGB input for relative pose estimation. Our analyses reveal its applicability to varying numbers of robots, diverse connection types, and distinct trajectory overlaps.
In this work, the relative extrinsic camera parameters are optimized. Optimizing the intrinsic camera parameters becomes particularly advantageous when dealing with various camera settings on different robots. Furthermore, the base NeRF model for each robot has limitations in handling unbounded scenes. A more advanced NeRF model, specially designed for unbounded scenes, can be seamlessly integrated into this project. Finally, in this study, we assume that the local camera poses are known but with different origins. By jointly optimizing NeRF and local camera poses, it becomes possible to eliminate the requirement for prior camera pose information.
VI Acknowledgments
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC).

### v2

@@ IV-C Trajectory Overlap Analysis
In this experiment, the sensitivity of the convergence of the relative poses with respect to the overlaps between the views of the robots is analyzed. For this experiment, two robots are considered. There are no common images between the robots, and the overlaps are only in the trajectories. Fig. 9 shows the top-down view of the overlaps, ranging from 5% to 40%, i.e. [5, 10, 20, 30, 40]%. The overlaps are determined using $x$ and $y$ coordinates of camera poses, assuming the poses are known in a coordinate frame. To divide the views among $N$ robots, first, the $x-y$ plane containing the desired object (the Chair, in this case) is divided into $N$ sectors, centered on the object. This will generate a 0% overlap for $N$ robots, which then is extended to increase the overlap, as shown in Fig. 9.
For two robots, once the overlaps are determined, the poses of robot $R^{2}$ are manipulated to create a relative displacement with respect to $R^{1}$, by translating them for 3m along each axis and rotating them for 30 around each axis. Then optimization is performed, assuming an initial estimate for the relative pose with zero for each translation component and each rotation.
@@ IV-D Waymo dataset - Unbounded Scenes
The San Francisco Mission Bay dataset (Waymo) [8] captures 12,000 images over 100 seconds along a 1.08 km route, using 12 cameras mounted on a car in an urban environment with dynamic objects and reflective surfaces. A 286m data segment with 233 images from a single roof-mounted camera providing a full surround view was selected for this experiment (Fig. 11). The sequence was divided into six segments to resemble a multi-agent setup. Initial relative poses are offset by 1m from actual values. For larger offsets, the algorithm fails due to lack of overlap.
Fig. 11 shows the rendering result for the baseline and Di-NeRF. All robots can render a view that they have never seen directly. The average values of PSNR and SSIM are 25.60 and 0.831 over 6 robots and 25.39 and 0.848 for the baseline. Table V shows detailed results. We divided the Waymo dataset into segments to benchmark Di-NeRF against Block-NeRF [8]. Table V shows the results. The objective of this experiment is not to surpass Block-NeRF [8], which is centralized. The goal is to achieve comparable results between distributed and centralized systems.
@@
Figure 10: A sequence of a Waymo trajectory is divided into six segments, i.e. robots, and used to evaluate Di-NeRF.
@@
Figure 11: Rendered images from Di-NeRF on the Waymo dataset, all six robots demonstrated the capability to render similar images with quality comparable to that achieved through centralized processing.
IV-E Convergence of the Relative Poses
In this section, convergence analysis with different initial relative pose values for the Barn sequence is provided. For simplicity, two robots are considered. The actual relative translation is 3m in all directions (i.e. x, y, and z), and the average initial values for three translations start from 0m to 2m (off by -1m to 1m), with increments of 0.02m. Similarly, the rotation ground truth is 45, and the initial value ranges from 0 to 90 (off by -45 to 45), with increments of 0.9. The translation error averages (0.86, 0.76)cm, and the rotation error (0.32, 0.38), as shown by the blue curve in Fig. 14. Experimentally, the algorithm converges even with initial values far from the ground truth.
Figure 12: (Left) Trajectories of two robots from the Barn sequence showing no overlaps in views. Rendered views after Di-NeRF training: (Top Right) with unknown relative pose; (Bottom Right) with a known relative pose.
Figure 13: Average values for relative rotations and translations for $R^{2}$ and $R^{3}$, in two cases: $R^{1}$ doesn’t fail (red and blue lines) and $R^{1}$ fails (green and purple lines) after 150 iterations. When $R^{1}$ fails, the global coordinate is transferred to $R^{2}$, and $R^{3}$ is optimizing for relative pose with respect to $R^{2}$.
IV-F No Overlap Setup
Even when there is no direct path overlap, scenes may be viewed from varying angles, allowing for relative pose calculation. For scenarios with no view overlaps, we have included an experiment using the Barn sequence. Fig. 13 depicts the trajectories for two robots with results. Di-NeRF struggles to compute relative poses with no view overlaps. However, knowing the relative poses, Di-NeRF merges the two models in a distributed fashion despite no absolute overlap.
IV-G Reference Robot Failure
If the robot chosen as the global coordinate frame fails or does not communicate, the global frame is switched to another robot. Fig. 13 shows an experiment where after 150 iterations, the reference robot is changed. Di-NeRF successfully optimizes the relative pose with respect to a new origin.
IV-H Technical Details of the Experiments
In this paper, three datasets are used to evaluate Di-NeRF: the synthetic dataset [2] (800$\times$800 pixels), the Tanks and Temples [37] (1920$\times$1080 pixels), and the Waymo dataset [8] (1217$\times$1096 pixels). For all datasets, the batch size is 2048, the learning rate is 0.01, the weight for depth loss $\gamma$ in Eq. (4) is 0.01 and the value of $\rho$ in Eq. (9) is 0.001. $\rho$ is the weight for the quadratic term and is the step size in the gradient ascent for the dual variable optimization. The number of steps before a communication round is 200. There are 5 communication rounds in all experiments.
NeRF Network Configuration: The network $\theta$ has 8 layers with 256 channels each. The hashmap size is $2^{15}$ for synthetic sequences and $2^{19}$ for the Barn and Waymo sequences. Table VII shows the training times for a fully connected graph. The memory required for a NeRF model is shown in Table VII. A robot needs 40.7 MB to exchange models, which can be reduced by compressing them when bandwidth is limited.
Table IV: Di-NeRF and baseline (Bsln.) performances for the Mission Bay sequence for 6 robots. $R\#1$ is the reference frame.
R#1 R#2 R#3 R#4 R#5 R#6 Bsln.
@@ IV-H Technical Details of the Experiments
SSIM $\uparrow$ 0.826 0.854 0.815 0.819 0.838 0.839 0.848
$\delta R^{\circ}\downarrow$ 2.89 2.86 2.81 2.91 2.79 -
$\delta t$(cm)$\downarrow$ - 4.2 4.0 4.5 4.2 4.7 -
Table V: Di-NeRF and Block-NeRF, on the Mission Bay sequence.
Block-NeRF (centralized) Di-NeRF (distributed)
No. Blocks PSNR SSIM No. Agents PSNR SSIM
2 25.65 0.871 2 25.22 0.866
8 24.56 0.892 8 25.41 0.854
16 27.1 0.903 16 25.34 0.861
Table VI: Di-NeRF training time comparison (sec/itr/robot).
Dataset #Robots Di-NeRF Di-NeRF;known $T_{i}^{g}$ Baseline
Synthetic Dataset 2 4.27 0.98 0.94
Tank and Temple 5 4.78 0.98 0.93
Mission Bay 6 4.73 0.97 0.95
Table VII: Images and NeRF size with the required bandwidth
Sequence Seq. Size (MB) Model Size (MB) BW per Link
Chair 31.04 4.07 40.7
Lego 42.9 4.07 40.7
Hotdog 40.9 4.07 40.7
Barn 481.5 64.07 640.7
Waymo 428.2 64.07 640.7
Figure 14: Impact of initial translation and rotation values on relative pose convergence (the Barn sequence) and performance differences between two approaches in optimizing the Di-NeRF model. The initial translation deviation from ground truth ranges from -1m to 1m, and for rotation from -45 to 45. The first approach (shown in blue) integrates monocular depth estimation into the relative pose optimization, while the second approach (depicted in red) does not incorporate monocular depth estimation.
IV-I Ablation Study
We conducted an ablation study to examine the effects of incorporating monocular depth estimation into relative pose optimization. The findings are detailed in Fig 14. While joint optimization of NeRF and relative pose often faces convergence issues with inaccurate initial position error estimates, the use of monocular depth estimation significantly mitigates this problem by correcting depth distortions.
@@ V Conclusion and Future Work
We introduced Di-NeRF, the first fully distributed NeRF used for multi-robot systems and 3D scenes. We demonstrate its versatility across synthetic and real datasets. Di-NeRF harnesses the advantages of NeRF in distributed learning, relying on RGB input for relative pose optimization. Our analyses reveal its applicability to varying numbers of robots, diverse connection types, and distinct trajectory overlaps.
In this work, the relative camera extrinsics are optimized. Optimizing the intrinsics is useful when cameras on the robots have different settings. Also, the base NeRF model for each robot has limitations in unbounded scenes. A model, designed for unbounded scenes, can be integrated into this project. Finally, we assume the local camera poses are known but with different origins. By jointly optimizing NeRF and local camera poses, one can eliminate this requirement. Also, by optimizing the architecture of NeRF, Di-NeRF can be efficient in bandwidth use compared to directly sending RGB images. This demonstrates Di-NeRF’s advantage in distributed systems, where communication is localized, reducing bandwidth needs unlike in centralized systems that require comprehensive data transmission. The core strength of Di-NeRF lies in facilitating collaborative mapping in challenging environments, emphasizing the distributed nature of learning and system scalability where broad communication may not be feasible.